# SoK: Honeypots & LLMs, More Than the Sum of Their Parts?

### v1

@@ Abstract
The advent of Large Language Models (LLMs) promised to resolve the long-standing paradox in honeypot design: achieving high-fidelity deception with low operational risk. However, despite a flurry of research since late 2022, progress has been incremental, and the field lacks a cohesive understanding of the emerging architectural patterns, core challenges, and evaluation paradigms. To fill this gap, this Systematization of Knowledge (SoK) paper provides the first comprehensive overview of this new domain. We survey and systematize three critical, intersecting research areas: first, we provide a taxonomy of honeypot detection vectors, structuring the core problems that LLM-based realism must solve; second, we synthesize the emerging literature on LLM-honeypots, identifying a canonical architecture and key evaluation trends; and third, we chart the evolutionary path of honeypot log analysis, from simple data reduction to automated intelligence generation. We synthesize these findings into a forward-looking research roadmap, arguing that the true potential of this technology lies in creating autonomous, self-improving deception systems to counter the emerging threat of intelligent, automated attackers.
@@ 1 Introduction
Cybersecurity honeypots—decoy systems designed to lure and analyze attackers—have long grappled with a fundamental design paradox. The spectrum of honeypots ranges from “low-interaction” systems, which are safe and easy to deploy but simplistic and easily detected, to “high-interaction” systems, which offer high-fidelity engagement at the cost of significant risk and management overhead. For decades, the central goal of honeypot research has been to resolve this tension: to achieve the high fidelity of a real system with the low risk and effort of a simulated one.
@@ 1.1 Related Works & Our Contributions
To situate our work, we conducted a review of recent surveys and Systematization of Knowledge (SoK) papers. Our search on Google Scholar for the terms ("survey" OR "systemization of knowledge" OR "SoK") AND "honeypot" AND ("Large Language Model" OR "LLM") from 2024 to the present yielded no prior SoKs on this specific topic.
@@ 1.2 Summarized Findings
Our systematization reveals several key findings regarding the current state of the art. Historic honeypot fingerprinting principles often still apply as even with LLM simulation, fundamental limitations will give these systems away to a reasonably intelligent adversary.
Architecturally, we identify an emerging canonical model for LLM-honeypots that incorporates components for filtering, state management, and generation, but we note a significant gap in open-source tooling. We find the field has developed a valuable trichotomy of evaluation criteria and methodologies: deceptiveness (human evaluation), accuracy (statistical), overall (real-world deployments). Yet there persists an evaluation paradox: while multiple methodologies exist, their utility is hampered by a real-world ”data desert” of low-sophistication traffic, making it difficult to measure progress against meaningful threats. This leads to our finding that the appropriate target for these systems is not skilled humans, whom they are unlikely to deceive, but the emerging threat of sophisticated automated agents. Finally, we find that while automated threat intelligence generation is now possible, it remains in a nascent, proof-of-concept stage, primarily blocked by the lack of large-scale, labeled training data.
Based on these findings, we propose a forward-looking research roadmap centered on the development of autonomous and self-improving deception systems. Architecturally, we highlight the need for open-source, modular frameworks and research into efficient, privacy-preserving models using techniques like LoRA. We advocate for the creation of a dynamic adversarial research ecosystem to solve the ”data desert” problem and enable meaningful evaluation against the true target: sophisticated, automated attackers. Finally, we outline a path toward true autonomy through the implementation of feedback loops, enabling systems to automatically generate intelligence for security operations centers and reconfigure themselves to adapt to the evolving threat landscape.
@@ 2.1 Network Cybersecurity Background
Cybersecurity is commonly defined as ensuring the confidentiality, integrity, and availability of data and networked computing assets. The practice of cybersecurity begins with strategically designing a network with a good security posture (e.g., protecting important assets via layers of defense with increasing restrictions, employing the principle of least privilege). This is complemented by a ”defense-in-depth” strategy, which uses many heterogeneous tools for prevention (e.g., firewalls, anti-malware), detection (e.g., signature-based or anomaly-based detectors), and management (e.g., vulnerability scanning) [9]. The overall effect is that well-defended networks enjoy protection against most previously seen attack patterns. This provides asymmetry, as attackers who use known methods have a decreased probability of success and are forced to develop or acquire new attack patterns, which is presumably more costly.
@@ 2.1.1 Cyber Attack Modeling Frameworks
Frameworks for understanding cyber attacks provide a common language to facilitate attack identification, detection, and threat intelligence sharing. The term TTP (Tactics, Techniques, and Procedures) is common in attack analysis. In this context, tactics are high-level adversarial objectives (e.g., Reconnaissance). Techniques are the methods used to achieve a tactic (e.g., Port Scanning). Finally, procedures are the specific commands or code used to implement a technique (e.g., a specific nmap command).
@@ 2.2 Honeypot Background
Honeypots are decoy computing systems designed to enhance security by gathering threat intelligence and occupying attacker resources. Consequently, honeypots have been a defensive tactic in cybersecurity since at least the 1980s [20, 21]. By convincing unknowing attackers to interact with them, honeypot systems allow for the direct observation of attackers’ actions, which in turn informs defensive mechanisms. This obviates the tedious process of finding malicious activity in a sea of benign system logs. Furthermore, honeypots are generally isolated environments designed so they can be attacked with no risk to production network assets.
@@ 2.2.1 Honeypot Goals & Levels of Simulation
With these goals in mind, honeypots mimic servers with varying degrees of realism and risk. They are conceptualized on a spectrum from low-interaction to high-interaction. Low-interaction honeypots are almost fully simulated with limited capabilities. Because they have little real connectivity, they pose minimal risk to the defender’s network, but they lack the realism to entice sophisticated attacker behavior. Medium-interaction honeypots have some real functionality but are restricted to limit risk. High-interaction honeypots are generally real systems, which require more resources to deploy and entail greater risk. Many templates for these honeypot types are open-sourced [30]. In general, lower interaction means faster deployment and less risk, but at the cost of being less convincing. Experts have stated that low-interaction honeypots will likely never convince a skilled human attacker [31], and our survey (Section 3.1) shows there are many ways to detect a simulated system. The vision driving the integration of LLMs into honeypot systems is to achieve the fidelity of a high-interaction honeypot with the speed, ease, and low risk of a lower-interaction system.
@@ 3 Related Fields
In this section we survey three areas, namely, honeypot fingerprinting methods, LLM shell honeypots, and honeypot data analysis methods. In each we organize the works into either categories or developing trends to admit a clearer understanding of the area.
@@ 3.1.1 Contents & Network Posture
Early detection methods focused on probing for environmental and network inconsistencies. The early work of Provos [42] created a honeypot network named Honeyd. Provos took care to camouflage the system’s fingerprint from scanning tools like NMAP, as well as adjusting outgoing packets (e.g., ISN, IP, and TCP characteristics) to maintain consistency. Subsequent research quickly found new inconsistencies with honeypots. Holz & Raynal showed that simple Bash commands or hardware details (e.g., MAC addresses) could reveal a simulated environment [43]. Defibaugh-Chavez et al. [44] illustrate many methods of honeypot detection, namely, exercising many different services (identifying ones that should exist but either do not or are unbelievable) and examining packet contents. These red flags have been reported for almost 20 years.
@@ 3.2.3 Architectures Advancement & Real-World Deployments
Building on a similar hybrid concept, Guan et al. [52] also considered LLMs simulating shell responses in a honeypot. Like Wang et al., Guan et al. instantiated a honeypot system with a filtering/routing step to prevent scanning activity from engaging the LLM. The filter leverages signatures to identify scanning scripts. Notably, Guan et al. provided statistics showing that a honeypot’s typical query rate exceeds API limits, proving that this DoS vulnerability is naturally exploited by the high rate of attacks. Their filtering logic keeps rates well within these limits. Guan et al. focus in part on prompting and provide guidance and tests for it. Finally, they provided statistics from “in-vitro” (laboratory) tests and an in-the-wild deployment using many different LLMs alongside a Cowrie-only honeypot. For laboratory tests, they constructed 50 attack sessions and tested each LLM-honeypot with a “vanilla” versus a chain-of-thought prompt, identifying the session length until an unbelievable output occurs. Across the board, chain-of-thought prompting provides a large boost. From a three-week internet deployment, they found that for effective attacks (those not deemed to be scanning), the average session length was 2.96 commands with Cowrie, which was statistically significantly smaller than an average of 5.83 commands from the best LLM-powered version.
@@ 3.3.3 From Patterns to Intent: Automated Mapping to MITRE ATT&CK
An early example is the XT-Pot framework from Ryandy et al. [92], which mapped observed attacker activity to ATT&CK techniques, presumably using manually defined rules to create ”soft signatures.” Gizzarelli [69] introduced an LLM-powered honeypot, SYNAPSE, which includes a machine learning classifier to automatically map attacker activity to the MITRE ATT&CK framework. When evaluated against real-world attack data, this mapping extension achieved 75% precision and 68% recall.
@@ 4.2.2 No Target Adversary
The pre-LLM honeypot era held the belief that low-interaction honeypots will never trick an intelligent human adversary [23], and despite the hype, our findings extend this belief to current LLM-driven honeypots. This stems from fundamental limitations that cannot be solved by realistic simulation; examples persist throughout honeypot history (see Section 3.1), such as the inability to convincingly simulate vim or htop outputs without custom architectures, and the even harder challenge of safely downloading and running an adversary’s file. These findings beg the question, ”If humans will not be tricked and bots have limited attack functionality, what is the target adversary of such systems?” Without a clear attacker model, targeted development of next-generation honeypots is not possible.
@@ 4.3 Automated Threat Intelligence
The most significant recent breakthrough in honeypot data analysis is the newfound ability to automate the labeling of attacker behavior with MITRE ATT&CK tactics, largely thanks to fine-tuned language models [94]. This marks the ”Semantic Leap” from processing raw data to generating structured intelligence. However, this research is still in a nascent stage. The current state-of-the-art, LogPrécis, is a proof-of-concept, limited by the lack of large-scale, well-labeled datasets. A clear need for real-world use is a sufficiently capable labeler, which depends on the creation of a larger and more attack-representative dataset.

### v3

@@ Abstract
The advent of Large Language Models (LLMs) promised to resolve the long-standing paradox in honeypot design, achieving high-fidelity deception with low operational risk. Through a flurry of research since late 2022, steady progress from ideation to prototype implementation is exhibited. Since late 2022, a flurry of research has demonstrated steady progress from ideation to prototype implementation. While promising, evaluations show only incremental progress in real-world deployments, and the field still lacks a cohesive understanding of the emerging architectural patterns, core challenges, and evaluation paradigms. To fill this gap, this Systematization of Knowledge (SoK) paper provides the first comprehensive overview and analysis of this new domain. We survey and systematize the field by focusing on three critical, intersecting research areas: first, we provide a taxonomy of honeypot detection vectors, structuring the core problems that LLM-based realism must solve; second, we synthesize the emerging literature on LLM-powered honeypots, identifying a canonical architecture and key evaluation trends; and third, we chart the evolutionary path of honeypot log analysis, from simple data reduction to automated intelligence generation. We synthesize these findings into a forward-looking research roadmap, arguing that the true potential of this technology lies in creating autonomous, self-improving deception systems to counter the emerging threat of intelligent, automated attackers.
@@ 1 Introduction
Cybersecurity honeypots—decoy systems designed to lure and analyze attackers—have long grappled with a fundamental design paradox. The spectrum of honeypots ranges from “low-interaction” systems, which are safe and easy to deploy but simplistic and easily detected, to “high-interaction” systems, which offer high-fidelity engagement at the cost of significant risk and management overhead. For decades, the central goal of honeypot research has been to resolve this tension: to achieve the high fidelity of a real system with the low risk and minimal effort of a simulated one.
@@ 1.1 Related Works & Our Contributions
To situate our work, we conducted a review of recent surveys and Systematization of Knowledge (SoK) papers. Our search on Google Scholar for the terms ("survey" OR "systematization of knowledge" OR "SoK") AND "honeypot" AND ("Large Language Model" OR "LLM") from 2024 to the present yielded no prior SoKs on this specific topic.
@@ 1.2 Summarized Findings
Our systematization reveals several key findings regarding the current state of the art. Historical honeypot fingerprinting principles often still apply. Even with LLM simulation, fundamental limitations will give these systems away to a reasonably intelligent adversary.
Architecturally, we identify an emerging canonical model for LLM-honeypots that incorporates components for filtering, state management, and generation, but we note a significant gap in open-source tooling. We find the field has developed a valuable trichotomy of evaluation criteria and methodologies: deceptiveness (human evaluation), accuracy (statistical), overall (real-world deployments). Yet there persists an evaluation paradox: while multiple methodologies exist, their utility is hampered by a real-world “data desert” of low-sophistication traffic, making it difficult to measure progress against meaningful threats. This leads to our finding that the appropriate target for these systems is not skilled humans, whom they are unlikely to deceive, but the emerging threat of sophisticated automated agents. Finally, we find that while automated threat intelligence generation is now possible, it remains in a nascent, proof-of-concept stage, primarily blocked by the lack of large-scale, labeled training data.
Based on these findings, we propose a forward-looking research roadmap centered on the development of autonomous and self-improving deception systems. Architecturally, we highlight the need for open-source, modular frameworks and research into efficient, privacy-preserving models using techniques like LoRA. We advocate for the creation of a dynamic adversarial research ecosystem to solve the “data desert” problem and enable meaningful evaluation against the true target: sophisticated, automated attackers. Finally, we outline a path toward true autonomy through the implementation of feedback loops, enabling systems to automatically generate intelligence for security operations centers (SOCs) and reconfigure themselves to adapt to the evolving threat landscape.
@@ 2.1 Network Cybersecurity Background
Cybersecurity is commonly defined as ensuring the confidentiality, integrity, and availability of data and networked computing assets. The practice of cybersecurity begins with strategically designing a network with a good security posture (e.g., protecting important assets via layers of defense with increasing restrictions, employing the principle of least privilege). This is complemented by a “defense-in-depth” strategy, which uses many heterogeneous tools for prevention (e.g., firewalls, anti-malware), detection (e.g., signature-based or anomaly-based detectors), and management (e.g., vulnerability scanning) [9]. The overall effect is that well-defended networks enjoy protection against most previously seen attack patterns. This provides asymmetry, as attackers who use known methods have a decreased probability of success and are forced to develop or acquire new attack patterns, which is presumably more costly.
@@ 2.1.1 Cyber Attack Modeling Frameworks
Frameworks for understanding cyber attacks provide a common language to facilitate attack identification, detection, and threat intelligence sharing. The term TTPs (Tactics, Techniques, and Procedures) is common in attack analysis. In this context, tactics are high-level adversarial objectives (e.g., Reconnaissance). Techniques are the methods used to achieve a tactic (e.g., Port Scanning). Finally, procedures are the specific commands or code used to implement a technique (e.g., a specific nmap command).
@@ 2.2 Honeypot Background
Honeypots are decoy computing systems designed to enhance security by gathering threat intelligence and occupying attacker resources. Consequently, honeypots have been a defensive tactic in cybersecurity since at least the 1980s [20, 21]. By convincing unwitting attackers to interact with them, honeypot systems allow for the direct observation of attackers’ actions, which in turn informs defensive mechanisms. This obviates the tedious process of finding malicious activity in a sea of benign system logs. Furthermore, honeypots are generally isolated environments designed so they can be attacked with no risk to production network assets.
@@ 2.2.1 Honeypot Goals & Levels of Simulation
With these goals in mind, honeypots mimic servers with varying degrees of realism and risk. They are conceptualized on a spectrum from low-interaction to high-interaction. Low-interaction honeypots are almost fully simulated with limited capabilities. Because they have little real connectivity, they pose minimal risk to the defender’s network, but they lack the realism to entice sophisticated attack behavior. Medium-interaction honeypots have some real functionality but are restricted to limit risk. High-interaction honeypots are generally real systems, which require more resources to deploy and entail greater risk. Many templates for these honeypot types are open-sourced [30]. In general, lower interaction means faster deployment and less risk, but at the cost of being less convincing. Experts have stated that low-interaction honeypots will likely never convince a skilled human attacker [31], and our survey (Section 3.1) shows there are many ways to detect a simulated system. The vision driving the integration of LLMs into honeypot systems is to achieve the fidelity of a high-interaction honeypot with the speed, ease, and low risk of a lower-interaction system.
@@ 3 Related Fields
In this section we survey three areas, namely, honeypot fingerprinting methods, LLM shell honeypots, and honeypot data analysis methods. In each, we organize the works into either categories or developing trends to admit a clearer understanding of the area.
@@ 3.1.1 Contents & Network Posture
Early detection methods focused on probing for environmental and network inconsistencies. The early work of Provos [42] created a honeypot network named Honeyd. Provos took care to camouflage the system’s fingerprint from scanning tools like NMAP, as well as adjusting outgoing packets (e.g., ISN, IP, and TCP characteristics) to maintain consistency. Subsequent research quickly found new inconsistencies with honeypots. Holz & Raynal showed that simple Bash commands or hardware details (e.g., MAC addresses) could reveal a simulated environment [43]. Defibaugh-Chavez et al. [44] illustrate many methods of honeypot detection, namely, exercising many different services (identifying ones that should exist but either do not or are unconvincing) and examining packet contents. These red flags have been reported for almost 20 years.
@@ 3.2.3 Architectures Advancement & Real-World Deployments
Building on a similar hybrid concept, Guan et al. [52] also considered LLMs simulating shell responses in a honeypot. Like Wang et al., Guan et al. instantiated a honeypot system with a filtering/routing step to prevent scanning activity from engaging the LLM. The filter leverages signatures to identify scanning scripts. Notably, Guan et al. provided statistics showing that a honeypot’s typical query rate exceeds API limits, proving that this DoS vulnerability is naturally exploited by the high rate of attacks. Their filtering logic keeps rates well within these limits. Guan et al. focus in part on prompting and provide guidance and tests for it. Finally, they provided statistics from “in-vitro” (laboratory) tests and an in-the-wild deployment using many different LLMs alongside a Cowrie-only honeypot. For laboratory tests, they constructed 50 attack sessions and tested each LLM-honeypot with a “vanilla” versus a chain-of-thought (CoT) prompt, identifying the session length until an unbelievable output occurs. Across the board, chain-of-thought prompting provides a large boost. From a three-week internet deployment, they found that for effective attacks (those not deemed to be scanning), the average session length was 2.96 commands with Cowrie, which was statistically significantly smaller than an average of 5.83 commands from the best LLM-powered version.
@@ 3.3.3 From Patterns to Intent: Automated Mapping to MITRE ATT&CK
An early example is the XT-Pot framework from Ryandy et al. [92], which mapped observed attacker activity to ATT&CK techniques, presumably using manually defined rules to create “soft signatures.” Gizzarelli [69] introduced an LLM-powered honeypot, SYNAPSE, which includes a machine learning classifier to automatically map attacker activity to the MITRE ATT&CK framework. When evaluated against real-world attack data, this mapping extension achieved 75% precision and 68% recall.
@@ 4.2.2 No Target Adversary
The pre-LLM honeypot era held the belief that low-interaction honeypots will never trick an intelligent human adversary [23], and despite the hype, our findings extend this belief to current LLM-driven honeypots. This stems from fundamental limitations that cannot be solved by realistic simulation; examples persist throughout honeypot history (see Section 3.1), such as the inability to convincingly simulate vim or htop outputs without custom architectures, and the even harder challenge of safely downloading and running an adversary’s file. These findings beg the question, “If humans will not be tricked and bots have limited attack functionality, what is the target adversary of such systems?” Without a clear attacker model, targeted development of next-generation honeypots is not possible.
@@ 4.3 Automated Threat Intelligence
The most significant recent breakthrough in honeypot data analysis is the newfound ability to automate the labeling of attacker behavior with MITRE ATT&CK tactics, largely thanks to fine-tuned language models [94]. This marks the “Semantic Leap” from processing raw data to generating structured intelligence. However, this research is still in a nascent stage. The current state-of-the-art, LogPrécis, is a proof-of-concept, limited by the lack of large-scale, well-labeled datasets. A clear need for real-world use is a sufficiently capable labeler, which depends on the creation of a larger and more attack-representative dataset.
@@ 5.2.3 Potential Architectural Directions
5.2.4 REST/JSON API Proposal
While this work focuses on shell-based honeypots, the extension of LLM simulation to REST/JSON APIs represents a promising avenue for future research. As the fundamental backend for mobile and web applications, these APIs constitute a critical attack surface exposed to external threats. The generative capabilities of LLMs are uniquely suited to mimicking complex APIs by producing dynamic, realistic-looking responses that are necessary to sustain attacker engagement. This high-fidelity simulation creates an opportunity to capture sophisticated probing for business logic flaws and other vulnerabilities. We anticipate that such honeypots would observe a broad spectrum of attacks, ranging from credential stuffing and data scraping to command injection attempts aimed at the underlying infrastructure fronted by the HTTP server.